I am in calculus online and just need some refreshers on how to rationalize denominators such as
4 divided by 3+ square root of 7 ( I came to an answer of -4(3-sqrt7)divided by 63)
or
2 divided by the square root of 10 (my answer -2 square root10)
I do know that you have to multiply the top and bottom to get the square root on top. Thank you!

Found 2 solutions by solver91311, John10:
Answer by solver91311(24713)



Your process is correct, and your numerator is correct, although it needs to be simplified, but your denominator is way out of whack. You have the right idea, you just need to work on your arithmetic.

Here's a hint: The product of a pair of conjugates is the difference of two squares, that is to say:



Your denominator should come out to be

Just use the distributive property to simplify the numerator.


John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by John10(297)

Hi,
To rationalize the denominator, you want to make sure that there is no radical sign in the denominator. In order to do so, you can multiply the denominator AND numerator by the conjugate of the denominator.
For example: 4/(3 + SQRT(7))
To rationalize, you multiply the denominator and numerator by (3 - SQRT(7)
[4(3 - sqrt(7)] / [(3 + sqrt(7))(3 - sqrt(7)]
= (12 - 4sqrt(7))/(9 - 7)
= (12 - 4sqrt(7)/ 2
= 6 - 2sqrt(7)
You can do the same for other problem.
By the ways, you always use the difference of two squares
(a - b)(a + b) = a^2 - b^2
Message me if you need extra assistance for your math class.
John